# Estimating treatment effects with competing intercurrent events in randomized controlled trials

### v1

@@ Abstract
The analysis of randomized controlled trials is often complicated by intercurrent events—events that occur after treatment initiation and may impact outcome assessment. These events may lead to patients discontinuing their assigned treatment or dropping out of the trial entirely. In an analysis of data from two recent immunology trials, we categorize intercurrent events into two broad types: those unrelated to treatment (e.g., withdrawal from the study due to external factors like pandemics or relocation) and those related to treatment (e.g., adverse events or lack of efficacy). We adopt distinct strategies to handle each type, aiming to target a clinically more relevant estimand. For treatment-related intercurrent events, they often meaningfully describe the patient’s outcome, we employ a composite variable strategy, where we attribute an outcome value that reflects the lack of treatment success. For treatment-unrelated intercurrent events, we adopt a hypothetical strategy that assumes these event times are conditionally independent of the outcome, given treatment and covariates, and envisions a scenario in which the intercurrent events do not occur. We establish the nonparametric identification and semiparametric estimation theory for the causal estimand and introduce doubly robust estimators. We illustrate our methods through the re-analysis of two randomized trials on baricitinib for Systemic Lupus Erythematosus. We classify intercurrent events, apply four estimators, and compare our approach with common ad-hoc methods, highlighting the robustness and practical implications of our framework.
Keywords: Causal inference; Censoring; Estimand; Post-treatment variable; Potential outcome
@@ 1 Intercurrent events in randomized controlled trials
Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy. However, after treatment initiation, various events—referred to as intercurrent events (ICEs)—can arise, impacting the interpretation or availability of the outcome of interest, and posing significant challenges to the analysis of RCTs. These ICEs can include treatment discontinuation due to adverse events, lack of efficacy, or study withdrawal due to patient relocation. Addressing these ICEs carefully is crucial to ensure the validity and reliability of the causal conclusions drawn from RCTs.
Given the complexities introduced by ICEs, clear guidance on how to handle them is crucial for drawing valid causal conclusions. The International Conference on Harmonization (ICH) has established global standards for drug development to align practices among pharmaceutical companies and regulatory agencies. Recognizing the importance of well-defined estimands in the presence of ICEs, the ICH introduced the E9 (R1) Addendum (ICH E9 , 2019, R1), a key resource that provides updated statistical principles to support clear and consistent approaches to analyzing and interpreting RCT results. The guidance has become essential for companies in drug research and development, and various strategies have been proposed and utilized to address ICEs following the guidance (Qu et al.,, 2021; Han and Zhou,, 2023).
@@
One straightforward approach to address the complications from the ICEs is the intention-to-treat (ITT) analysis. The ITT strategy includes all participants in their originally assigned groups and uses the observed outcome value, regardless of whether they completed the treatment or adhered to the study protocol. This strategy, also known as the treatment policy strategy in ICH E9 , 2019 (R1), is relevant for policy decisions and widely accepted. However, it fails to address ICEs that are terminal events, such as death, since these events directly impact the existence of the outcome variable.
To address the limitation of ITT analyses and approach the issue from a different perspective, the composite variable strategy modifies the causal parameter of interest from the commonly used average causal effect to a comparison between two composite potential outcomes across the treatment arms. The composite potential outcome is defined as the endpoint of interest if no ICE occurs before its measurement, and a pre-specified, clinically meaningful value indicates treatment failure otherwise. The composite variable strategy is widely used in practice across different types of endpoints: binary endpoints (e.g., the non-responder imputation approach), ordinal or continuous endpoints (e.g., Rosenbaum,, 2006), and time-to-event endpoints (e.g., progression-free survival). In addition to the treatment policy strategy and the composite outcome strategy, three additional strategies for addressing ICEs were provided in ICH E9 , 2019 (R1). The hypothetical strategy considers treatment effects under a counterfactual scenario where ICEs do not occur or occur differently. The remaining two strategies–the while-on-treatment strategy and the principal stratification strategy–are less commonly used in practice. ICH E9 , 2019 (R1) recommends using different strategies based on the specific type of ICEs involved. However, very few papers have formally explored how to combine different strategies and rigorously addressed the pitfalls of doing so naively (Lipkovich et al.,, 2020; Qu et al.,, 2021).
1.1 A phase-3 immunology trial
To illustrate our setup and notation, we first introduce a real phase-3 immunology trial (Morand et al.,, 2023; Petri et al.,, 2023), which we will revisit in Section 6. This trial investigates the causal effect of baricitinib versus placebo on Systemic Lupus Erythematosus. Patients in the RCT were assigned either to a baricitinib treatment group or to a placebo control group.
@@ 1.1 A phase-3 immunology trial
The primary endpoint of interest is an immune response index measured 52 weeks after treatment initiation. Ideally, we would compare this outcome across the two groups at week 52. However, 429 out of 1535 (27.95%) patients experienced ICEs between treatment initiation and the outcome measurement, leading to missingness of the outcome. Figure 1 shows the types and proportions of these ICEs. During the 52 weeks, various ICEs happened. For instance, some patients discontinued the assigned treatment due to adverse events or lack of efficacy, some had protocol deviation and were excluded from the study, some of them withdrew from the study for unknown reasons, and some were lost to follow-up.
Figure 1: Pie chart showing the ICE types and proportions
@@ 1.2 Our proposal and contribution
We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are likely correlated with both treatment and outcomes. Figure 1 illustrates these classifications: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and blue stripped charts represent a mix of both types. For instance, patients withdrew from the study for various reasons–some due to relocation or the pandemic, which can be considered as unrelated to the treatment, while others withdrew due to concerns about potential side effects, which are likely to be treatment-related. When patients did not disclose a detailed reason, we adopt a conservative approach and classify their ICEs as treatment-related.
Both types of ICEs may occur between treatment initiation and outcome measurement. Figure 2 depicts the timeline and three scenarios in our motivating example. Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to the treatment, patient 2 dropped out of the study due to an adverse event, and patient 3 completed the study through the endpoint at week 52, when the outcome of interest was measured.
Figure 2: Illustration of the motivating immunology trial example
Both types of ICEs lead to missing primary endpoints measured at a landmark time because the outcome cannot be observed or is not even well-defined when an ICE occurs. For the first type of ICE, we apply the hypothetical strategy and assume the ICE event happens at random conditional on the observed covariates. It considers the hypothetical scenario in which the ICE did not occur. For the second category, we utilize the composite variable strategy, incorporating the ICE itself as indicative of treatment failure, for which we provide a formal definition and discussion in Section 3.
A main challenge in combining the two strategies and our key insight is that the two types of ICEs can compete with each other, meaning that patients discontinue due to the first reason that occurs. Use study withdrawal due to relocation and adverse events as examples to represent the two types of ICEs, respectively. For example, suppose a patient discontinues due to relocation like patient 1 in Figure 2. In that case, it remains unknown whether an adverse event would have occurred before the final endpoint measurement in the counterfactual scenario had the withdrawal due to relocation ICE not occurred. It is important to carefully deal with the issue of competing ICEs, which, to the best of our knowledge, has never been pointed out in the literature. We consider this feature and provide the nonparametric identification of our causal parameter of interest under two assumptions: (1) treatment assignment satisfies the ignorability and overlap assumptions, and (2) the treatment-unrelated ICE time is independent of the potential outcomes and the potential values of the treatment-related ICE time, given observed covariates. We present two identification formulas relying on different sets of nuisance models. We then propose two basic estimators based on these two identification formulas and a third estimator that combines them in a similar way to the construction of the augmented inverse probability weighting estimator to achieve double robustness.
The combined estimator can be further improved in terms of both robustness and asymptotic efficiency. To construct a more principled estimator that simultaneously achieves robustness and semiparametric efficiency, we derive the efficient influence function (EIF) thus the semiparametric efficiency bound of our causal parameter. We propose an EIF-based estimator and study its asymptotic properties. We show the EIF estimator is doubly robust, meaning its consistency does not require all nuisance parameters to be consistently estimated—only one of the two sets of nuisance models needs to be correctly specified. Moreover, the EIF estimator achieves the semiparametric efficiency bound under certain conditions on nuisance estimation.
@@
The remaining of the paper proceeds as follows. In Section 2, we provide a motivating example, introduce the basic setup of our research question, describe the composite variable strategy, define our causal parameter of interest, and illustrate the identification challenge. In Section 3, we present the key identification assumptions, provide the nonparametric identification of our causal parameter, and construct three basic estimators. In Section 4, we derive the EIF for our causal parameter, propose a doubly robust and asymptotically efficient estimator based on the EIF, and study its asymptotic properties. In Section 5, we run Monte Carlo simulations to evaluate the finite sample performance of our proposed estimation procedures. In Section 6, we implement the estimators to re-analyze two randomized controlled trials that study the causal effect of baricitinib treatment on systemic lupus erythematosus. In Section 7, we discuss two future research directions.
@@
We use the following notation. Let $\lVert r\rVert_{2}=\{\int r(o)^{2}\textnormal{d}P(o)\}^{1/2}$ denote the $L_{2}(P)$ norm where $P(\cdot)$ denotes the distribution of the observed data $O=o$. For the survival and censoring functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm. We write $b_{n}=O_{P}(a_{n})$ if $b_{n}/a_{n}$ is bounded in probability and $b_{n}=o_{P}(a_{n})$ if $b_{n}/a_{n}$ converges to 00 in probability.
2 Basic setup
2.1 Setup and notation
@@
In this subsection, we introduce our formal setup and mathematical notation. Let $A$ denote the binary treatment indicator, with $A=1$ for patients in the treatment group, and $A=0$ for those in the control group. The primary outcome, denoted as $Y$, is measured at a pre-specified time point $k$. In our motivating example, $Y$ is the immune response index and $k$ is 52 weeks. We adopt the potential outcomes framework (Neyman,, 1923; Rubin,, 1974), where $Y(a)$ denotes the potential outcome under treatment assignment $A=a$ for $a=0,1$, and $Y=Y(A)=AY(1)+(1-A)Y(0)$ is the observed outcome.
@@
As discussed, two types of ICEs may occur between the treatment initiation and the measurement of the outcome: treatment-unrelated ICEs, assumed independent of the potential outcomes conditional on covariates, and treatment-related ICEs, which likely correlate with the endpoints. To formally define the two types of ICEs, let $C$ and $T$ denote time to treatment-unrelated and time-to-treatment-related ICEs, respectively. Both are post-treatment variables, thus we use $C(a)$ and $T(a)$ to denote their potential values under treatment $a$, for $a=0,1$, with observed values $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
2.2 Causal parameter of interest and challenge in identification
@@
To address treatment-related ICEs, we define a composite potential outcome as $Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$ for $a=0,1$ where the superscript stands for “composite” and define our parameter of interest as the mean contrast, i.e.,
@@
This target parameter is widely used for binary outcomes to capture success or failure, where treatment-related ICEs, such as adverse events, are treated as failure (ICH E9 , 2019, R1). In our motivating example, $Y(a)$ measures whether a patient is an immune responder at week 52 after taking the treatment $A=a$, which is unobserved for patients experiencing a treatment-related ICE such as an adverse event or lack of efficacy. Such patients are defined as non-responders under the composite variable strategy.
@@
The composite variable approach extends beyond binary outcomes to accommodate scenarios where a clinically significant value can denote treatment failure. More specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are measured on the ordinal and continuous scale, if there is a predefined clinically meaningful value $v$ that can signal treatment failure, then a composite outcome under treatment arm $A=a$ may be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, in which case the causal parameter is defined as in (1) with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$. Rosenbaum, (2006) discussed similar causal parameters when the ICE is death and the outcome of interest is some measure of life quality after a certain period.
Identifying $\tau$ in the presence of both ICE types is challenging. For clarity, in the remainder of this subsection, we focus on withdrawal due to relocation and adverse events to represent the treatment-unrelated and treatment-related ICEs, respectively. The definition of the causal parameter $\tau$ in (1) essentially treats $Y^{\textup{cp}}(a)$ as the relevant outcome, assigning 0 when adverse events occur, i.e., when $T(a)<k$. If adverse events were the only reason for missingness in the outcome, we could identify $\tau$ using the difference in means of the observed composite outcomes $Y^{\textup{cp}}=Y^{\textup{cp}}(A)=Y1(T>k)$, i.e.,
$\tau$ $=$ $E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}.$
@@
However, this formula is infeasible to learn from observed data because treatment-unrelated ICEs, such as withdrawal due to relocation, could happen before treatment-related ICEs thus censoring $T(a)$. Figure 2 shows how withdrawal due to relocation in patient 1 censored their adverse event time and the outcome, creating missingness in the composite outcome $Y1(T>k)$ for such patients. The key challenge is that the two types of ICEs compete with each other, and patients discontinue due to the first ICE that occurs. Table 1 summarizes observed ICE types and composite outcomes, highlighting that censoring complicates the identification of $\tau$. When there is no ICE, we observe the outcome $Y$, which equals the composite outcome because $T>k$. When we observe an adverse event, the composite outcome is by definition 0. When we observe a withdrawal due to relocation, both the outcome of interest and the time-to-adverse event are missing, and so is the composite outcome. We address the challenge and present identification results in the following section.
@@
Table 1: Summary of the observed ICE types and outcome. $C$ is the time to withdrawal due to relocation (WR), $T$ is the time to adverse event (AE), $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken.
@@
no AE/WR $C\wedge T>k$ $Y$ 1 $Y$
AE $C\wedge k>T$ 0 0
WR $T\wedge k>C$
3 Identification and three basic estimators

### v2

@@ Abstract
The analysis of randomized controlled trials is often complicated by intercurrent events (ICEs) – events that occur after treatment initiation and affect either the interpretation or existence of outcome measurements. Examples include treatment discontinuation or the use of additional medications. In two recent clinical trials for systemic lupus erythematosus with complications of ICEs, we classify the ICEs into two broad categories: treatment-related (e.g., treatment discontinuation due to adverse events or lack of efficacy) and treatment-unrelated (e.g., treatment discontinuation due to external factors such as pandemics or relocation). To define a clinically meaningful estimand, we adopt tailored strategies for each category of ICEs. For treatment-related ICEs, which are often informative about a patient’s outcome, we use the composite variable strategy that assigns an outcome value indicative of treatment failure. For treatment-unrelated ICEs, we apply the hypothetical strategy, assuming their timing is conditionally independent of the outcome given treatment and baseline covariates, and hypothesizing a scenario in which such events do not occur. A central yet previously overlooked challenge is the presence of competing ICEs, where the first ICE censors all subsequent ones. Despite its ubiquity in practice, this issue has not been explicitly recognized or addressed in previous data analyses due to the lack of rigorous statistical methodology. In this paper, we propose a principled framework to formulate the estimand, establish its nonparametric identification and semiparametric estimation theory, and introduce weighting, outcome regression, and doubly robust estimators. We apply our methods to analyze the two systemic lupus erythematosus trials, demonstrating the robustness and practical utility of the proposed framework.
Keywords: Causal inference; Clinical trial; International Council for Harmonization; Post-treatment variable; Potential outcomes
@@ 1 Intercurrent events in randomized controlled trials
Randomized controlled trials (RCTs) are considered the gold standard for evaluating treatment efficacy, primarily because randomization supports assumption-lean inference of the treatment effect. However, after treatment initiation, various events, referred to as intercurrent events (ICEs), can arise, impacting the interpretation or availability of outcome measurements and posing significant challenges to the analysis of RCTs. Examples of ICEs include treatment discontinuation due to adverse events or lack of efficacy, patient relocation, and the use of additional medications. Carefully accounting for ICEs is essential to ensure the validity and reliability of the causal conclusions drawn from RCTs.
Recognizing the critical need for clearly defined estimands in the presence of ICEs, the International Council for Harmonization (ICH) issued the E9(R1) Addendum (ICH E9 , 2019, R1). This addendum introduces a structured estimand framework for clinical trials to obtain precisely defined treatment effects that align with the clinical questions of interest. It outlines strategies for addressing ICEs during the formulation of the clinical question and emphasizes that careful specification of the treatment, population, and outcome variable often addresses many of the ICEs raised in discussions between sponsors and regulators. Since its release, the ICH E9(R1) Addendum has been widely discussed, increasingly adopted in clinical drug development, and has sparked substantial interest in statistical research (Qu et al.,, 2021; Kang et al.,, 2022; Ionan et al.,, 2023; Han and Zhou,, 2023; Olarte Parra et al.,, 2025).
@@
A widely accepted strategy is the treatment policy strategy (ICH E9 , 2019, R1), which includes all participants in their originally assigned groups and uses the observed outcome values, regardless of whether or not an ICE occurs. This approach aligns with the intention-to-treat principle and reflects the effect of a treatment policy in real clinical settings. However, it cannot address ICEs that are terminal events, such as death, because such events preclude the existence of the outcome variable.
The composite variable strategy is an alternative approach that is well-suited for handling ICEs that are informative of the patient’s outcome such as the terminal events. This strategy incorporates the occurrence of ICEs directly into the outcome definition. Specifically, it defines a composite outcome: if no ICE occurs, the outcome of interest is used as observed; if an ICE occurs, a pre-specified, clinically meaningful value—typically indicating treatment failure—is assigned. This strategy is widely used in practice across various types of outcomes, including binary outcomes (e.g., the non-responder imputation approach), ordinal or continuous outcomes (e.g., Rosenbaum,, 2006), and time-to-event outcomes (e.g., progression-free survival).
In addition to the treatment policy and composite variable strategies, ICH E9 , 2019 (R1) outlines three additional strategies for handling ICEs. The hypothetical strategy evaluates treatment effects under a hypothetical scenario in which the ICE would not occur. The while-on-treatment and principal stratification strategies are also described but are less commonly used in current practice due to their reliance on strong assumptions and the potential to introduce bias in treatment comparisons. ICH E9 , 2019 (R1) recommends using different strategies based on the specific type of ICEs involved. While some studies have explored the use of multiple strategies (Lipkovich et al.,, 2020; Qu et al.,, 2021), none have rigorously examined how to combine different strategies, nor have they rigorously addressed the potential pitfalls of doing so without careful consideration. Moreover, the challenges posed by competing ICEs remain unrecognized.
1.1 Two phase-3 trials in systemic lupus erythematosus
To rigorously examine the challenges posed by ICEs in clinical trials, we analyze data from two recent randomized studies for systemic lupus erythematosus (Morand et al.,, 2023; Petri et al.,, 2023), in which participants were randomly assigned to either the baricitinib treatment group or the placebo control group. These twin trials were designed to provide substantiated evidence on the causal treatment effect of baricitinib versus placebo.
@@ 1.1 Two phase-3 trials in systemic lupus erythematosus
The primary outcome is a response index measured at 52 weeks after treatment initiation. Ideally, this outcome would be compared directly between the two treatment groups at week 52. However, 429 out of 1,535 patients (27.95%) experienced ICEs during the follow-up period, resulting in unobserved outcome data. Figure 1 shows the types and proportions of these ICEs. During the 52-week period, various ICEs occurred: some patients discontinued treatment due to adverse events or lack of efficacy; others discontinued treatment due to study withdrawal for unspecified reasons or were lost to follow-up; and some were excluded from the study due to protocol deviations. We revisit this example in greater detail in Section 5.
Figure 1: Pie chart showing the ICE types and proportions. The yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and the blue striped charts represent a mix of both types.
@@ 1.2 Our proposal and contribution
We propose to classify ICEs into two broad types: (1) treatment-unrelated ICEs, such as treatment discontinuation due to relocation or COVID-19 lockdown, which are assumed to be independent of treatment efficacy conditional on the observed covariates; and (2) treatment-related ICEs, such as treatment discontinuation due to adverse events or lack of efficacy, use of rescue medication, and terminating events such as death, which are often informative about a patient’s outcome.
In our application studies, the two types of ICEs were classified based on a manual review of the detailed comments collected at the clinical sites. Figure 1 summarizes the main categories and illustrates our classification scheme: yellow solid charts represent treatment-related ICEs, the green solid chart represents a treatment-unrelated ICE, and the blue striped charts represent a mix of both types. For instance, some patients withdrew from the study due to external factors such as relocation or the COVID-19 pandemic, which were classified as treatment-unrelated. Others withdrew due to concerns about potential side effects, which were considered treatment-related. When the reason for withdrawal was not clearly stated, we adopted a conservative approach and classified the ICE as treatment-related. Both types of ICEs may occur during the follow-up period. Figure 2 depicts the timeline and three representative scenarios from our motivating example. Before week 52, patient 1 withdrew from the study due to relocation, which is arguably unrelated to treatment. Patient 2 dropped out due to an adverse event, reflecting a treatment-related ICE. Patient 3 completed the study through week 52, at which point the primary outcome was measured.
Figure 2: Illustration of three representative scenarios of intercurrent events in the immunology trial. The primary outcome is measured at week 52, and the solid lines indicate the observed follow-up period for each patient. Patient 1 experienced a treatment-unrelated ICE (represented by an arrow), Patient 2 experienced a treatment-related ICE (represented by a diamond), and Patient 3 had no ICE before week 52. This figure highlights the issue of competing ICEs: for instance, had Patient 1 not discontinued early and remained on treatment (illustrated by the gray dashed line), they might have experienced a treatment-related ICE later in the follow-up period.
Both types of ICEs lead to unmeasured primary outcomes at the pre-specified landmark time because the outcome cannot be observed or may not even be well-defined when a treatment-related ICE occurs. To ensure the estimand is clinically interpretable and relevant, we apply different strategies tailored to the nature of the ICE. For ICEs that are plausibly unrelated to treatment (e.g., treatment discontinuation due to relocation or administrative withdrawal), we apply the hypothetical strategy, which imagines a scenario where the ICE did not occur. This enables estimation of the treatment effect as if the patient had remained on treatment and in the study as planned, under the assumption that the ICE time is conditionally independent of the outcome given treatment and baseline covariates. This assumption is reasonable in settings where the ICE arises from external or administrative factors that are not influenced by post-randomization conditions. For example, if a patient relocates for personal reasons unrelated to their health status or treatment response, then, conditional on baseline characteristics and assigned treatment, the occurrence of this ICE can be viewed as independent of the patient’s outcome. In contrast, for treatment-related ICEs (e.g., treatment discontinuation due to adverse events or lack of efficacy), we apply the composite variable strategy, treating the ICE itself as an indication of treatment failure, as such events signify that the patient is unable to continue treatment. A formal definition and discussion of this strategy are provided in Section 3. Combining the two strategies enhances the interpretability of the estimand and ensures it better reflects clinically relevant questions, as no single strategy can adequately address both types of ICEs. Specifically, it would be inappropriate to classify a patient as a non-responder if they discontinue treatment due to relocation, and it is of limited clinical relevance to consider a hypothetical scenario in which adverse events do not occur. This approach is consistent with regulatory guidance, which recommends applying different strategies depending on the type of ICE involved (ICH E9 , 2019, R1; Kang et al.,, 2022).
However, a central yet previously overlooked challenge is the presence of competing ICEs, where the first ICE censors all subsequent ones. For example, if a patient discontinues treatment due to relocation, as illustrated by patient 1 in Figure 2, we only observe that no treatment-related ICE occurred before the discontinuation. However, it remains unknown whether an adverse event would have occurred before the final outcome measurement in the hypothetical scenario in which the patient had not relocated and had continued treatment as planned. Although such competing ICEs are ubiquitous in practice, this issue has not, to our knowledge, been explicitly recognized or addressed in prior methodological work.
The central thesis of this paper is to address the challenge of competing ICEs. We begin by developing a principled framework that clearly defines the estimand and establishes its nonparametric identification. Specifically, we derive two identification formulas, each relying on a different set of nuisance parameters. Building on these results, we propose two basic estimators corresponding to the two identification strategies. We then introduce an augmented estimator that combines the two, achieving double robustness. To further enhance both robustness and efficiency, we derive the efficient influence function (EIF) and construct an EIF-based estimator that attains the semiparametric efficiency bound under appropriate conditions. We apply our methods to the two systemic lupus erythematosus trials, both partially impacted by the COVID-19 pandemic, demonstrating the robustness and practical utility of the proposed framework.
1.3 Organization and notation
@@
The remainder of the paper is organized as follows. In Section 2, we present a motivating example, introduce the basic setup of our research question, define the causal parameter of interest, and highlight the identification challenge. In Section 3, we state the key identification assumptions, establish the nonparametric identification of the causal parameter, and construct three estimators. In Section 4, we derive the EIF, propose an estimator that is both doubly robust and asymptotically efficient, and examine its asymptotic properties. In Section 5, we apply the methods to re-analyze the two systemic lupus erythematosus trials. Finally, in Section 6, we conclude with a discussion of two directions for future research. The supplementary material includes a simulation study and all proofs. We also provide publicly available R code implementing all four proposed estimators.
@@
We use the following notation. Let $\lVert r\rVert_{2}=\{\int r(o)^{2}\textnormal{d}P(o)\}^{1/2}$ denote the $L_{2}(P)$ norm where $P(\cdot)$ denotes the distribution of the observed data $O=o$. For the survival functions, let $\lVert r\rVert_{2}=\{\iint r(t,o)^{2}\textnormal{d}P(o)\textnormal{d}t\}^{1/2}$ denote the $L_{2}(P)$ norm. We write $b_{n}=O_{P}(a_{n})$ if $b_{n}/a_{n}$ is bounded in probability and $b_{n}=o_{P}(a_{n})$ if $b_{n}/a_{n}$ converges to 00 in probability.
2 Setup and estimand
2.1 Setup
@@
Let $A$ denote the binary treatment indicator, where $A=1$ corresponds to assignment to the treatment group and $A=0$ to the control group. The primary outcome, denoted by $Y$, is measured at a pre-specified time point $k$. In our motivating example, $Y$ represents the response index, and $k$ corresponds to 52 weeks. We adopt the potential outcomes framework, in which $Y(a)$ denotes the potential outcome under treatment assignment $A=a$, for $a=0,1$. The observed outcome is then given by $Y=Y(A)=AY(1)+(1-A)Y(0)$.
@@
As discussed in Section 1, two types of ICEs may occur after treatment initiation: treatment-unrelated ICEs and treatment-related ICEs. To formally define these, let $C$ and $T$ denote the time to a treatment-unrelated ICE and the time to a treatment-related ICE, respectively. Since both are post-treatment variables, we use $C(a)$ and $T(a)$ to denote their potential values under treatment assignment $A=a$, for $a=0,1$. The observed times are then given by $C=C(A)=AC(1)+(1-A)C(0)$ and $T=T(A)=AT(1)+(1-A)T(0)$.
2.2 Causal estimand and challenges in identification
@@
To address treatment-related ICEs, we define a composite potential outcome as
$Y^{\textup{cp}}(a)=Y(a)1\{T(a)>k\}$
for $a=0,1$, where the superscript denotes “composite”. This construction defines the outcome as zero (i.e., failure) if a treatment-related ICE occurs before the outcome is measured. We then define our causal estimand as the mean contrast:
@@
This is the estimand under the composite variable strategy. It is widely used for binary outcomes to represent treatment success or failure, where treatment-related ICEs such as adverse events or lack of efficacy are treated as failures (ICH E9 , 2019, R1). In our motivating example, $Y(a)$ indicates whether a patient would be a responder at week 52 under treatment assignment $A=a$, had no ICE occurred. For patients who experience a treatment-related ICE before week 52, $Y(a)$ is unobserved. Under the composite variable strategy, these patients are defined as non-responders.
@@
The composite variable approach extends beyond binary outcomes and is applicable in settings where treatment failure can be defined using a clinically meaningful threshold. Specifically, for outcomes such as quality of life, chronic pain, physical functioning, and cognitive performance that are typically measured on an ordinal or continuous scale, if a predefined value $v$ may be used to indicate treatment failure, then a composite outcome under treatment assignment $A=a$ can be constructed as $Y(a)1\{T(a)>k\}+v1\{T(a)\leq k\}$, where the outcome retains its actual value if no treatment-related ICE occurs before time $k$, and takes the failure value $v$ otherwise. The corresponding causal estimand can then be defined as in (1), with $Y(a)$ replaced by $Y(a)-v$ for $a=0,1$. Rosenbaum, (2006) discussed a similar causal parameter in settings where the ICE is death and the outcome of interest is a measure of quality of life assessed after a fixed period.
Identifying $\tau$ in the presence of both types of ICEs presents a key methodological challenge. For clarity, in the remainder of this subsection, we use treatment discontinuation due to relocation and adverse events to represent treatment-unrelated and treatment-related ICEs, respectively. The causal estimand $\tau$ defined in (1) is based on the composite potential outcome $Y^{\textup{cp}}(a)$, which takes the value 00 if an adverse event occurs before the outcome measurement, i.e., when $T(a)<k$. If adverse event were the only type of ICE, and treatment is randomized, $\tau$ could be identified from observed data using the difference in means of the composite outcomes: $\tau=E\{Y1(T>k)\mid A=1\}-E\{Y1(T>k)\mid A=0\}$. However, this formula is infeasible in the presence of treatment-unrelated ICEs, which may occur before treatment-related ICEs and thus censor both $T(a)$ and $Y$. As illustrated in Figure 2, patient 1 discontinued due to relocation, which censored both the occurrence of a potential adverse event and the outcome, resulting in an unobserved composite outcome $Y{1}(T>k)$. Therefore, for patients who experience treatment-unrelated ICEs, the problem cannot be treated as standard censoring, and one cannot naïvely impute the outcome $Y$ under a hypothetical scenario where the treatment-unrelated ICE did not occur. This is because such patients may still have experienced a treatment-related ICE had they not discontinued due to relocation. In other words, a valid identification strategy must recover the expected value of the composite outcome $Y{1}(T>k)$, rather than the outcome $Y$ alone.
@@
Table 1 summarizes the observed ICE types and corresponding composite outcomes, highlighting that competing ICEs complicate the identification of $\tau$. When no ICE occurs, the outcome $Y$ is observed and equals the composite outcome since $T>k$. When an adverse event is observed, the composite outcome is, by definition, 0. However, when a patient discontinues treatment due to relocation, both the outcome $Y$ and the time to the adverse event $T$ are unobserved, and so is the composite outcome. We address this challenge and present formal identification results in the following section.
@@
Table 1: Summary of the observed ICE types and outcome. $C$ is the time to treatment discontinuation due to relocation, $T$ is the time to adverse event, $Y$ is the outcome of interest, and $k$ is the pre-specified time point when the measurement of $Y$ is taken. A question mark “?” indicates that the corresponding value is unobserved.
@@
no any type of ICE $C\wedge T>k$ $Y$ 1 $Y$
adverse event $C\wedge k>T$ 0 0
treatment discontinuation due to relocation $T\wedge k>C$
3 Nonparametric identification and basic estimators